Please, be neat and detailed. Explanations would be great. I need to understand it. Thank you.

Question 122 points total). Your sister is starting to plan for your nieces college expenses and has asked for your advice.
Answer #1
Age Year Deposits FV
6 0 5,000.00 10,060.98 =5000*(1.06)^12
7 1 6,500.00 12,338.94 =6500*(1.06)^11
8 2 8,000.00 14,326.78
9 3 9,500.00 16,050.05
10 4 11,000.00 17,532.33
11 5 12,500.00 18,795.38
12 6 14,000.00 19,859.27
13 7 15,500.00 20,742.50
14 8 17,000.00 21,462.11
15 9 18,500.00 22,033.80
16 10 20,000.00 22,472.00
17 11 21,500.00 22,790.00
18 12 23,000.00 23,000.00 =23000*(1.06)^0
Total 241,464.13

Initially the amount will be invested for 12 years, then 11 years, then 10 years and will therefore each subsequent deposit will be calculated as per the table given above.

The formula to calculate the withdrawals is PMT(0.06,5,241,464.13) = -57322.72
This can be shown as per the schedule below:
Initial end deposit = 241,464.13*1.06
End deposits = End balance*1.06

Age Year End Deposits Withdrawals End balance
19 13 255,951.98 -57,322.72 198,629.26
20 14 210,547.02 -57,322.72 153,224.30
21 15 162,417.76 -57,322.72 105,095.05
22 16 111,400.75 -57,322.72 54,078.03
23 17 57,322.72 -57,322.72 0.00